# Enhanced Optimization Strategies to Design an Underactuated Hand Exoskeleton

### v1

@@ II-C2 Big Bang-Big Crunch Algorithm (BBBC)
output : The most fitting solution $\textnormal{{P}}(1)$
@@
7                   $\textnormal{{P}}\leftarrow\textnormal{{bang}}(\textnormal{{cm}},i)$;
@@
III Design Process I: Single-Objective Optimization with Mechanical Constraints
U-HEx was designed with a linear actuator (Firgelli L12-P) with a maximum displacement of 50 mm [19]. In this study, we evaluated the impact of presenting the mechanical limitations of the minimum and maximum linear actuator displacements on U-HEx design by comparing the optimality of their retrieved solutions against the original SOOP [20].
III-A Optimization Problem (SOOP)
@@
maximize $\sqrt{\tau_{MCP}+\tau_{PIP}}$ (1)
where $\tau_{MCP}$ and $\tau_{PIP}$ are calculated based on the values of the decision variables (link lengths). The optimization problem is subject to the following seven inequality constraints (one constraint for each inequality):
  • $con_{1-4}$: U-HEx is connected to the user’s fingers through passive linear sliders ($c_{1}$ and $c_{2}$) whose movements should be limited by the user’s finger size [19], resulting in four constraints defined in Eqn. (2) and (3);
    $0\leq c_{1}\leq 35$ (2)
    $0\leq c_{2}\leq 45$ (3)
  • $con_{5-6}$: to guarantee a solid grasp and user’s safety, the torques $\tau_{MCP}$ and $\tau_{PIP}$ should have a ratio within a reasonable range, resulting in two constraints defined in Eqn. (4) – this range comes from the static analysis from the original U-HEx work [19]; and
@@
  • $con_{7}$: to ensure a natural range of motion for hand closing, the desired actuator displacement ($L_{x}$) must comply with the chosen actuator’s mechanical spectrum, resulting in the constraint defined in Eqn. (5) – this constraint was not considered in our previous work [20].
@@
Each of these constraints is a function of the decision variables (link lengths) and is calculated in Simulink, similarly to the objective function (see Fig. 2(a) [19]). Additional constraints to the problem are related to the decision variable bounds (i.e., the range of link lengths) reported in Table I.
III-B Experiment Outline (SOOP)
To find the optimal design of U-HEx, we designed an experiment with three factors as (i) the constraint to the desired actuator displacement (with and without $con_{7}$), (ii) the optimization algorithm (GA and BBBC), and (iii) the number of decision variables (six and nine link lengths).
When the number of link lengths is limited to six, the additional three links are kept constant as $\overline{BK}=37$, $\overline{CI}=16$, and $\overline{EJ}=37$. These values were selected empirically as detailed in the original work [19]. To ensure a valid comparison, we kept the algorithm parameters the same as our previous work [20], reported in Table II. Each algorithm was executed 10 times on a computer with 16 core 5.4 GHz CPU and 64 GB RAM.
TABLE II: Parameter settings in SOOP for EAs
(Design Process I)
@@ III-B Experiment Outline (SOOP)
Selection Type (GA) Binary Tournament [26]
Crossover Type (GA) blx-$\alpha$ ($\alpha$ = 0.5) [27]
@@
Mutation Type (GA) Polynomial [28]
@@
Survival Strategy (GA) Elitist ($\mu+\lambda$ schema) [29]
Crunch Method (BBBC) Best Fit [30]
Constraint Handling (GA, BBBC) Deb’s Method [31]
III-C Evaluation Metrics (SOOP)
@@ III-C1 Optimality
III-C2 Convergence Time (CT)
CT indicates the time needed for algorithms to find the optimal solution, defined in Eqn (6).
@@ III-C2 Convergence Time (CT)
where the overall runtime (RT) indicates the overall time taken by the algorithm and the number of generations (NG) is 50 as indicated in Table II. The generation of convergence (GC) is the observed generation in which the algorithm converged to a solution – specifically if the variance in the value of the retrieved solution did not change in the last 20 iterations with a margin of 0.05 Nm.
III-D Results (SOOP)
TABLE III: Numerical Results of Design Process I
with 6 and 9 Link Lengths (LL), with and without
Actuation Displacement Constraint ($con_{7}$: $L_{x}\leq 50$)
@@ III-D Results (SOOP)
Figure 3: Force transmission (optimality) and convergence time comparisons between experiments (with a different number of variable link lengths): with the constraint to the maximum actuator displacement ($con_{7}=L_{x}\leq 50$) and without [20]. Plots report the mean and the standard error.
We analyzed optimality and convergence time using a three-way Repeated Measure Analysis of Variance (RM-ANOVA). Fig. 3 shows the mean and standard error of the collected data, as reported in Table III; whereas Table IV summarizes the ANOVA results.
III-D1 Results on Optimality
Fig. 3 shows that adding the actuation constraint ($con_{7}$) to the optimization problem significantly lowered the force transmission. The drop between with and without $con_{7}$ becomes higher with nine decision variables compared to six. We further analyzed the trends between optimization strategies and decision variables only with $con_{7}$ through a post-hoc analysis, as depicted by a zoomed-in version in Fig. 4. We found that the optimality results are statistically significantly higher with nine decision variables ($F(1,36)=1047621.937,p<0.001,\eta^{2}=1.000$) and while using BBBC ($F(1,36)=74.239,p<0.001,\eta^{2}=0.673$). In addition, we observed their interactions to be statistically significant ($F(1,36)=38.156,p<0.001,\eta^{2}=0.515$).
III-D2 Results on Convergence Time
@@
III-E Discussion (SOOP)
This experiment aimed to observe the impact of physical constraints on the optimization problem while using different EAs. We specifically added a constraint to the desired actuator displacement, which is given by the actuator chosen from the market (Firgelli L12-P), for a natural hand movement. We used two EAs (GA and BBBC) and compared our results to our previous study [20]. We hypothesized ($H_{1}$) that adding the additional constraint would lower the optimality (i.e., force transmission), regardless of the optimization technique. Our results in Fig. 3 show that including the actuation constraint lowers the force transmission while using both optimization methods and both decision variable settings (six and nine link lengths), as expected – therefore, $H_{1}$ holds true.
Furthermore, Fig. 3 shows that increasing the number of decision variables from six to nine significantly increased the optimality without the actuation constraint, but not with it. This might indicate that all the solutions leading to high values of optimality require actuator displacements higher than 50 mm (i.e., unfeasible solutions), which will be investigated in the following design process with MOOP (see Sec. IV). In addition to a decrease in optimality, we observe that adding the actuation constraint resulted in statistically higher convergence times for the optimization algorithms. We speculate that higher convergence times might be due to the change in the search space and the computational toll of evaluating one more constraint, supported by the evidence in the literature [31].
@@ III-E Discussion (SOOP)
TABLE IV: Statistical results for SOOP
in terms of Optimality and Convergence Time
@@
We also analyzed the differences between the two EAs and the number of decision variables when the actuation constraint was implemented separately since the optimality results are primarily influenced by the optimization problem without the actuation constraint. Supported by further statistical analyses, Fig. 4 shows that increasing the number of decision variables provides a statistical increase in force transmission – similar to our previous study [20].
@@
Interestingly, we observed that BBBC offers statistically higher optimality results than GA while working with six and nine decision variables independently – even though previously we found no difference between them [20]. We speculate that this was caused by the GA being stuck in a local maxima, which was significantly lower than the global maxima. This is also supported by observing the convergence time: while BBBC reached the current optimal solution [^1] We remind readers that EC methods retrieve approximate optimal solutions rather than the exact one. relatively fast, GA continued the search for a significantly longer time. This is also supported by the fact that the convergence time of GA did not change as much as BBBC when the number of decision variables was increased from six to nine. As mentioned in the state-of-the-art, BBBC was proposed to address some of the biggest disadvantages of GAs, such as premature convergence, convergence speed, and execution time [23].
@@
IV Design Process II: Multi-Objective Optimization with Mechanical Constraints
@@
The design process in SOOP (Sec. III) showed the importance of considering the actuator displacement. Introducing a constraint to the desired actuator displacement (Eqn. 5) to perform natural hand movements reduced the force transmission of the exoskeleton but allowed it to be equipped with smaller actuators that can fit on top of the hand while still exerting a sufficient amount of force to assist users in opening and closing their fingers. In the meantime, users’ safety was ensured by balancing the torques applied on both finger joints through two inequality constraints to bound the ratio between joint torques to be within a specific range as specified in Eqn. (4). Even though these constraints allowed us to have a design within safe and feasible boundaries, they do not ensure reaching an optimal one about these factors.
@@
Based on these considerations, we re-defined the optimization problem to include two further objectives that were previously treated as constraints (i.e., $con_{5-6}$ and $con_{7}$). We believe that turning constraints into additional objectives can help find better solutions by allowing more flexibility and balancing multiple requirements effectively, resulting in a set of different designs with different properties.
IV-A Background on Multi-Objective EAs (MOEAs)
@@
MOOP involves optimizing two or more objectives simultaneously. These objectives usually conflict with each other (i.e., optimizing one might worsen the second, and so on), and it is not possible to retrieve a single solution that fully optimizes all the objectives simultaneously. Therefore, optimality in MOOP is defined by the concept of pairwise dominance: a solution $x_{1}$ dominates a solution $x_{2}$ if $x_{1}$ is not worse than $x_{2}$ in all objectives and $x_{1}$ is strictly better than $x_{2}$ in at least one objective [32]. When solutions do not dominate each other, then they are considered as optimal. This results in a set of trade-off solutions, each with different degrees of gain and losses to the objectives. This set is referred to as the Pareto-Optimal Front, or just Pareto Front.
@@
EAs are great candidates for tackling MOOPs and looking for a set of solutions due to their population-based nature, allowing them to carry on different alternatives at a time [32]. These are usually referred to as Multi-Objectives Evolutionary Algorithms (MOEAs). MOEAs aim to retrieve an optimal set (named the Pareto set) with the following properties: (i) convergence, or proximity to the true Pareto Front (which is unknown for engineering problems), and (ii) diversity, ensuring that the solutions are well-spaced so that the Pareto Front is uniformly covered. These properties are satisfied by different types of Multi-Objective Survival Strategies by removing weak solutions from the population. MOEAs usually generate a number of offspring larger than the set population size (useful to promote space exploration); however, this creates a surplus in the population. Survival strategies deal with this surplus by promoting non-dominated isolated solutions (elites) and removing the ones located in crowded regions to promote diversity. For this study, we will explore two of the most well-known strategies in the state-of-the-art.

### v2

@@ 2.2.2 Big Bang-Big Crunch Algorithm (BBBC)
output : The most fitting solution $\textnormal{{P}}\left(1\right)$
@@
7                   $\textnormal{{P}}\leftarrow\textnormal{{bang}}\left(\textnormal{{cm}},i\right)$;
@@
3 Design Process I: Single-Objective Optimization with Mechanical Constraints
U-HEx was designed with a linear actuator (Firgelli L12-P) with a maximum displacement of 50 mm [sarac2017design]. In this study, we evaluated the impact of presenting the mechanical limitations of the minimum and maximum linear actuator displacements on U-HEx design by comparing the optimality of their retrieved solutions against the original SOOP [akbas2024impact].
3.1 Optimization Problem (SOOP)
@@
maximize $\sqrt{\tau_{MCP}\text{ + }\tau_{PIP}}$ (1)
where $\tau_{MCP}$ and $\tau_{PIP}$ are calculated based on the values of the decision variables (link lengths). The optimization problem is subject to the following seven inequality constraints:
  • 1.
    $con_{1-4}$: U-HEx is connected to the user’s fingers through passive linear sliders ($c_{1}$ and $c_{2}$) whose movements should be limited by the finger size [sarac2017design], as Eqn. (2);
    $0\leq c_{1}\leq 35;\quad 0\leq c_{2}\leq 45$ (2)
  • 2.
    $con_{5-6}$: to guarantee a solid grasp and user’s safety, the torques $\tau_{MCP}$ and $\tau_{PIP}$ should have a ratio within a reasonable range [sarac2017design], as Eqn. (3); and
@@
  • 3.
    $con_{7}$: to ensure a natural range of motion for the hand, the desired actuator displacement ($L_{x}$) must comply with the chosen actuator’s mechanical spectrum, as Eqn. (4) – this constraint was not considered in our previous work [akbas2024impact].
@@
Each of these constraints is a function of the decision variables (link lengths) and is calculated in Simulink, similarly to the objective function (see Fig. 3(a) [sarac2017design]). Additional constraints to the problem are related to the decision variable bounds (i.e., the range of link lengths) reported in Table 2.
3.2 Experiment Outline (SOOP)
We designed an experiment with three factors as (i) the constraint to the desired actuator displacement (with and without $con_{7}$), (ii) the optimization algorithm (GA and BBBC), and (iii) the number of decision variables (six and nine link lengths). When the number of link lengths is limited to six, the additional three links are kept constant as $\overline{BK}=37$, $\overline{CI}=16$, and $\overline{EJ}=37$. These values were selected empirically as detailed in the original work [sarac2017design]. To ensure a valid comparison, we kept the algorithm parameters the same as our previous work [akbas2024impact], reported in Table 3. Each algorithm was executed 10 times on a computer with 16 core 5.4 GHz CPU and 64 GB RAM.
Table 3: Parameter Settings in SOOP for EAs
@@ 3.2 Experiment Outline (SOOP)
Selection Type (GA) Binary Tournament [goldberg1991comparative]
Crossover Type (GA) blx-$\alpha$ ($\alpha$ = 0.5) [eshelman1993real]
@@
Mutation Type (GA) Polynomial [deb1996combined]
@@
Survival Strategy (GA) Elitist ($\mu+\lambda$ schema) [beyer2002evolution]
Crunch Method (BBBC) Best Fit [gencc2010big]
Constraint Handling (GA, BBBC) Deb’s Method [coello2002theoretical]
3.3 Evaluation Metrics (SOOP)
@@ 3.3.1 Optimality
3.3.2 Convergence Time (CT)
CT indicates the time needed for algorithms to find the optimal solution, defined in Eqn. (5).
@@ 3.3.2 Convergence Time (CT)
where the overall runtime (RT) indicates the overall time taken by the algorithm and the number of generations (NG) is 50 as indicated in Table 3. The generation of convergence (GC) is the observed generation in which the algorithm converged to a solution – specifically if the variance in the value of the retrieved solution did not change in the last 20 iterations with a margin of 0.05 Nm.
3.4 Results (SOOP)
We analyzed optimality and convergence time using a three-way Repeated Measure Analysis of Variance (RM-ANOVA). Fig. 4 shows the mean and standard error of the collected data, as reported in Table 4; whereas Table 5 summarizes the ANOVA results.
3.4.1 Results on Optimality
Fig. 4 shows that implementing the actuation constraint ($con_{7}$) significantly lowered the force transmission. The drop between with and without $con_{7}$ becomes higher with nine decision variables compared to six. We further analyzed the trends between optimization strategies and decision variables only with $con_{7}$ through a post-hoc analysis, as depicted by a zoomed-in version in Fig. 5. We found that the optimality results are statistically significantly higher with nine decision variables ($F(1,36)=1047621.937,p<0.001,\eta^{2}=1.000$) and while using BBBC ($F(1,36)=74.239,p<0.001,\eta^{2}=0.673$). In addition, we observed their interactions to be statistically significant ($F(1,36)=38.156,p<0.001,\eta^{2}=0.515$).
Figure 4: Force transmission (optimality) and convergence time comparisons between experiments (with a different number of variable link lengths): with the constraint to the maximum actuator displacement ($con_{7}=L_{x}\leq 50$) and without [akbas2024impact]. Plots report the mean and the standard error.
Table 4: Numerical Results of Design Process I (SOOP) with 6 and 9 Link Lengths (LL), with and without Actuation Displacement Constraint ($con_{7}$: $L_{x}\leq 50$)
@@ 3.4.1 Results on Optimality
3.4.2 Results on Convergence Time
@@
3.5 Discussion (SOOP)
We investigated the impact of constraints based on the desired actuator displacement (limitations of the actuator chosen from the market, Firgelli L12-P) on the optimization problem using different EAs. We used two EAs (GA and BBBC) and compared our results to our previous study [akbas2024impact]. Our results (Fig. 4) show that including the actuation constraint lowers the optimality (i.e., force transmission) while using both optimization methods and decision variable settings (six and nine link lengths), as expected – therefore, $H_{1}$ holds true.
Fig. 4 also shows that increasing the number of decision variables significantly did not increase the optimality when the actuator constraint was introduced. This might indicate that all the solutions leading to high values of optimality require actuator displacements higher than 50 mm (i.e., unfeasible solutions), which will be investigated in the following design process with MOOP (see Sec. 4). In addition to a decrease in optimality, we observed that the actuation constraint resulted in statistically higher convergence times for the optimization algorithms. We speculate that higher convergence times might be due to the change in the search space and the computational toll of evaluating one more constraint [coello2002theoretical].
Table 5: Statistical Results of Design Process I (SOOP)
@@ 3.5 Discussion (SOOP)
We also analyzed the differences between the two EAs and the number of decision variables when the actuation constraint was implemented separately since the optimality results were primarily influenced by the optimization problem without the actuation constraint. Fig. 5 shows that increasing the number of decision variables provides a statistical increase in force transmission – similar to our previous study [akbas2024impact].
@@
Interestingly, we observed that BBBC offers statistically higher optimality results than GA while working with six and nine decision variables independently – even though previously we found no difference between them [akbas2024impact]. We speculate that this was caused by the GA being stuck in a local maxima instead of a global one. This is also supported by observing the convergence time: while BBBC reached the current optimal solution [^1] We remind readers that EC methods retrieve approximate optimal solutions rather than the exact one. relatively fast, GA continued the search for a significantly longer time. Thus, the convergence time of GA did not change as much as BBBC when the number of decision variables was increased. As mentioned in the state-of-the-art, BBBC was proposed to address some of the biggest disadvantages of GAs, such as premature convergence, convergence speed, and execution time [erol2006new].
@@
4 Design Process II: Multi-Objective Optimization with Mechanical Constraints
@@
The design process in SOOP (Sec. 3) showed the importance of considering the actuator displacement as a constraint. Introducing a constraint to the desired actuator displacement (Eqn. 4) reduced the force transmission of the exoskeleton but allowed it to be equipped with smaller actuators that can fit on top of the hand while still exerting a sufficient amount of force to assist users in opening and closing their fingers. In the meantime, users’ safety was ensured by balancing the torques applied on both finger joints, as detailed in Eqn. (3).
@@
Even though these constraints allowed us to have a design within safe and feasible boundaries, they do not ensure reaching an optimal one about these factors. We re-defined the optimization problem to include two further objectives that were previously treated as constraints (i.e., $con_{5-6}$ and $con_{7}$). We believe that turning constraints into additional objectives can help find better solutions by allowing more flexibility and balancing multiple requirements effectively, resulting in a set of different designs with different properties.
4.1 Background on Multi-Objective EAs (MOEAs)
@@
MOOP involves optimizing two or more objectives; they usually conflict with each other (i.e., optimizing one might worsen the second), and it is not possible to retrieve a single solution that fully optimizes all the objectives simultaneously. Therefore, optimality in MOOP is defined by the concept of pairwise dominance: a solution $x_{1}$ dominates a solution $x_{2}$ if $x_{1}$ is not worse than $x_{2}$ in all objectives and $x_{1}$ is strictly better than $x_{2}$ in at least one objective [deb2001multi]. When solutions do not dominate each other, they are considered as optimal. This results in a set of trade-off solutions (i.e., Pareto Front), each with different degrees of gain and losses to the objectives.
@@
EAs are great candidates for tackling MOOPs and looking for a set of solutions due to their population-based nature, allowing them to carry on different alternatives at a time [deb2001multi]. Multi-Objectives Evolutionary Algorithms (MOEAs) aim to retrieve an optimal set (named the Pareto set) with the following properties: (i) convergence, or proximity to the true Pareto Front (which is unknown for engineering problems), and (ii) diversity, ensuring that the solutions are well-spaced so that the Pareto Front is uniformly covered. These properties are satisfied by different types of Multi-Objective Survival Strategies by removing weak solutions from the population. MOEAs usually generate a number of offspring larger than the set population size (to promote space exploration); however, this creates a surplus in the population. Survival strategies deal with this surplus by promoting non-dominated isolated solutions (elites) and removing the ones located in crowded regions to promote diversity. For this study, we will explore two of the most well-known strategies in the state-of-the-art.